# Underage Detection through a Multi-Task and MultiAge Approach for Screening Minors in Unconstrained Imagery

### v1

@@ Abstract
Accurate automatic screening of minors in unconstrained images demands models that are robust to distribution shift and resilient to the children under-representation in publicly available data. To overcome these issues, we propose a multi-task architecture with dedicated under/over-age discrimination tasks based on a frozen FaRL vision-language backbone joined with a compact two-layer MLP that shares features across one age-regression head and four binary under-age heads for age thresholds of 12, 15, 18, and 21 years, focusing on the legally critical age range. To address the severe class imbalance, we introduce an $\alpha$-reweighted focal-style loss and age-balanced mini-batch sampling, which equalizes twelve age bins during stochastic optimization. Further improvement is achieved with an age gap that removes edge cases from the loss.
Moreover, we set a rigorous evaluation by proposing the Overall Under-Age Benchmark, with 303k cleaned training images and 110k test images, defining both the “ASORES-39k” restricted overall test, which removes the noisiest domains, and the age estimation wild shifts test “ASWIFT-20k” of 20k-images, stressing extreme pose ($>$45°), expression, and low image quality to emulate real-world shifts.
Trained on the cleaned overall set with resampling and age gap, our multiage model “F” lowers the root-mean-square-error on the ASORES-39k restricted test from 5.733 (age-only baseline) to 5.656 years and lifts under-18 detection from F2 score of 0.801 to 0.857 at 1% false-adult rate. Under the domain shift to the wild data of ASWIFT-20k, the same configuration nearly sustains 0.99 recall while boosting F2 from 0.742 to 0.833 with respect to the age-only baseline, demonstrating strong generalization under distribution shift. For the under-12 and under-15 tasks, the respective boosts in F2 are from 0.666 to 0.955 and from 0.689 to 0.916, respectively.
@@
Age Estimation , Non-adult Detection , Distribution Shift, Class imbalance
@@
[…]
organization=Department of Electrical, Systems and Automation Engineering, Universidad de León, city=León, country=Spain
@@ 1 Introduction
Facial age estimation is the task of estimating a person’s age, given an image of the face [1]. As such, age estimation is a regression problem, but it is often interpreted and implemented as a classification problem, either by rounding the age to integer values (in years) or by defining a few larger age groups, as, for example, in [2]. The binary classification problem of only two age groups, below and above an age threshold, is known as age verification (is a person older than a given age level?), or, with opposite polarity, as non-adult detection or underage detection (is a person younger than a given age level?).
@@
Age estimation and its variants in form of age classification have become increasingly important across a wide range of applications. These include age-restricted services, targeted advertising, social media moderation, and, most critically, the protection of minors. Conceivable applications also involve equipping devices such as gambling machines, cigarette vending machines, or age-restricted smartphone applications with automated facial age-verification systems to enforce legal age limits. In particular, non-adult detection is essential in efforts to enforce age-related regulations in digital environments and to combat the distribution of Child Sexual Exploitation Material (CSEM). A possible pipeline for detecting CSEM hidden in huge amounts of data is a two-step process, consisting of detection of sexual activity [3] and detection of underage subjects (the focus of this contribution).
@@
While significant progress has been made in facial age estimation, most existing solutions are designed for controlled settings and tend to perform well when image conditions are ideal. However, real-world scenarios (“in the wild”) present substantial challenges, including variations in head pose, facial expressions, lighting conditions, image resolution, and frequent occlusions [4]. These factors severely impact the reliability of the model and often degrade its performance. These challenges become even more pronounced when we attempt to distinguish children, adolescents, and adults since this task is inherently more complex within these critical age ranges. This is precisely where accurate detection becomes more necessary for legal and safety considerations.
Furthermore, publicly available facial datasets, like MORPH [5], AFAD [6] or AgeDB [7], have a skewed distribution with a significant under-representation of children and adolescents. This imbalance leads to models with lower performance for younger age groups. Consequently, the effectiveness of underage detection systems is compromised in critical applications, such as CSEM investigations.
@@
In this work, we address the challenges of robust facial age estimation and non-adult detection by proposing a comprehensive framework for unconstrained environments. The main contributions of this study are as follows:
  • ASWIFT-20k, a novel wild test benchmark, explicitly designed to evaluate the robustness of the model under diverse and challenging imaging conditions, including variations in pose, facial expression, illumination, and image quality. In addition, we implement a systematic data cleaning process to reduce label noise and improve training reliability.
@@
  • ASORES-39k, an Overall Under-Age Benchmark with 303k cleaned training images and 110k test images created to generate age estimation and underage detection models with accuracy while excluding examples with noisy or unreliable labels.
@@
  • An age-balanced sampling strategy during training to address the over-representation of adult subjects in age-labeled datasets. This ensures equal representation of minors, enhancing the model ability to generalize across underrepresented age groups.
  • A novel $\alpha$-reweighted focal loss to mitigate class imbalance effects. This approach prioritizes difficult and minority-class samples, improving discriminative performance, particularly in underage detection tasks.
The proposed methodology aims to improve facial age estimation, particularly for distinguishing between prepubescent, pubescent, and adult individuals. Based on our experience of working with law enforcement agencies (LEAs) on previous projects, and considering the same age for males and females, we established the relevant ages to be 12 (prepubescent), 15 (pubescent) and 18 or 21 (adult), respectively. The focus is on developing robust underage detection systems that work well in varied real-world conditions, such as diverse poses, facial expressions, and poor image quality. In addition, it addresses key machine-learning challenges, such as class imbalance and domain shift.
The rest of the paper is organized as follows. Section 2 reviews related work in facial age estimation and non-adult detection. Sections 3 and 4 detail the proposed methodology, first the dataset construction, and then the model architecture and training procedures. Section 5 presents experimental results and discusses performance across different benchmarks. Finally, Section 6 summarizes the main conclusions.
@@ 2 Related Work
Facial age estimation history
During aging, the proportions of the face, the shape of facial features, skin textures (wrinkles, liver spots), and tone change. Early works on facial age estimation sought to model such features explicitly. Lanitis et al. [8] developed a statistical face model based on a principal component analysis of facial landmark coordinates and the isolation of the aging axis in PCA space. Han et al. [9] localized certain facial regions in the face image where they obtained biologically inspired features (BIF), which were hierarchical, then used in an approach of a binary decision tree of four age groups and four support-vector regressors, one for each group. Later approaches employed deep learning and convolutional neural networks (CNNs) [10], trained in an end-to-end fashion without the need for landmarks or hand-crafted features [11, 12]. Model sizes range from a few million parameters [11] to over 100M parameters in the case of the VGG-16 architecture used by Rothe et al. [12]. SSR-Net [13] is another CNN-based age-regression model, which uses a stage-wise strategy that allows reducing the size (less than 100k parameters) and the computational effort significantly. Li et al. [14] developed an adaptive label distribution learning, exploiting the ordinal character of the age classes of age estimation phrased as classification problem.
@@
Recently, Paplhám and Franc [15] compared state-of-the-art age estimation algorithms and provided an age-estimation benchmark, including train/validation/test splits for several facial age-estimation datasets. Considering the mean absolute error, averaged over the test part of the respective split, they found that the key to good age-estimation accuracy is a large pre-trained backbone, rather than the details of the loss function. The proposed model of Paplhám and Franc [15] combines the FaRL image encoder [16] with a small multilayer perceptron (MLP) for age estimation. FaRL is a general facial-representation model based on the general-purpose CLIP model [17], pre-trained with contrastive loss to learn the high-level semantic meaning from image-text pairs for a general facial representation. With 87M parameters, the image encoder of FaRL, a 12-layer vision transformer [18], is large compared to early CNNs and “slim” models like SSR-Net, but still smaller than the VGG-16 of Rothe et al. [12].
Facial non-adult detection history
@@
Meanwhile, other works have focused on the binary discrimination problem of non-adult detection. Castrillón-Santana et al. [19] composed a dataset of unrestricted images from a collection of datasets and compared local descriptors and CNNs, i.e., end-to-end learned features, for non-adult detection. They found that CNNs slightly outperformed local descriptors, but that score fusion among the two approaches could lead to even better accuracy. Anda et al. [20] have developed the DeepUAge age-estimation model, trained with the VisAGe dataset for accurate age estimation of underage subjects. Chaves et al. [21] have investigated facial age estimation of minors when the eye region is occluded, a situation that may occur in CSEM. Gangwar et al. [22] employs CNNs with an attention mechanism and metric learning for pornographic content detection and age-group classification and combines these into a CSEM detection pipeline.
3 Overall Under-Age Benchmark and ASWIFT-20K wild test
@@
This section will introduce the dataset and test benchmarks proposed: the Overall Underage Benchmark (Section 3.1.1), together with the Restricted Overall Test and the ASWIFT-20K (Section 3.2.2), designed explicitly for evaluating model robustness to distribution shift.
3.1 Overall Under-Age Benchmark
3.1.1 Raw composition and demographics
@@
For real-world applications of facial age estimation and underage detection (facial age verification), it is important that the age algorithm is reliable for all kinds of data; generalization and robustness with respect to distribution shifts (imaging conditions, demographics of the subjects, poses, facial expression) are crucial. Therefore, we merge several benchmarks, in the train/validation/test splits as provided by Paplhám and Franc [15], into an overall benchmark. The joint training set consists of data from the publicly available datasets AFAD [6], AgeDB [7], CLAP2016 [23], the Cross-Age Celebrity Dataset (CACD2000) [24], Morph Album 2 [5], and UTKFace [25]. These datasets cover different aspects of the data distribution in terms of age, image quality, poses, and facial expression. Figure 1 depicts some samples of the datasets used.
@@
Figure 1: Sample face patches from the different age-estimation benchmarks, as processed by the model during training, i.e., with the random flipping, scaling, and cropping described in Section 4.
@@
AFAD contains Asian subjects, the image resolution is rather low, and the face crops are tight, which leads to padding in the face patches. Morph Album 2 contains mugshot images, i.e. mostly frontal poses with neutral to negative facial expressions and controlled illumination. AgeDB, CACD2000, CLAPD2016, and UTKFace are “in-the-wild” datasets with diversity in pose, illumination, and expression. AgeDB also contains monochrome photographs. Of these six datasets, the CACD2000 dataset [24] stands out in the sense that it has noisy age annotations and was actually not intended for age estimation. Therefore, we investigate two variants of the overall benchmark, with and without the CACD2000 dataset.
@@
Figure 2 shows the age distribution of the overall training set, broken down into source datasets. Note that images of subjects below 15 years are only weakly represented, mostly from the datasets UTKFace, CLAP2016, and AgeDB.
Figure 2: The age distribution of the training data (cumulative histograms). (LEFT) Number of images for each source dataset (colors) and for the merged corpus (black). The steep rise of CACD2000 (green) and AFAD (blue) beyond 15 years, together with the almost flat curves of AgeDB, UTKFace, and CLAP2016 below that point, quantifies the severe under-representation of minors. (RIGHT): same curves for the whole age range normalized by the respective dataset sizes, which might reveal complementary age coverage: AFAD is concentrated in the 15 – 35 bracket, CACD2000 in 20 – 45, whereas UTKFace and CLAP2016 provide the bulk of the sub-15 data. The aggregate (black) displays the natural, highly skewed age distribution that motivates the age-balanced resampling scheme described in Section 4.2.
@@
We employ the data splits provided by Paplhám and Franc [15] and use the respective holdout parts of the training data for testing. Further, we use the FG-NET dataset [8] for testing only, as well as the Dartmouth Database of Children’s Faces [26] and the CASIA-Face-Africa dataset collected by the Chinese Academy of Science’s Institute of Automation (CASIA) [27].
3.1.2 Label-noise cleaning protocol
@@
We used a preliminary model for cleaning the data by inspecting the largest (apparent) training and validation errors. [^1] Largest (apparent) training and validation errors selected for manual inspection: (i) worst over-age classifications: images with labeled age $\geq 18$, select the lowest over-18 score, $0.5\%$ ($1738$ images); (ii) worst under-age classifications: images labeled with age $<18$, select the highest over-18 scores, $2\%$ ($293$ images); (iii) worst age underestimations: estimation error $\leq-16$ and estimate $<$24 ($1006$ images); (iv) worst age over-estimations: estimation error $\geq 9$) and labeled age $<18$ ($990$ images). The mentioned images were manually classified into the categories “bad age”, “doubtful”, “fine” and “bad image”. To avoid bias in favor of our own algorithm, doubtful cases were preserved, and only images with obviously wrong age labels and bad images (showing non-faces or cartoons) were removed. In another iteration, we also cleaned the test set (with more tolerant settings because, in general, test errors are larger than training errors). Altogether, 2777 out of 303075 images were removed from the training set, 238 out of 60220 from validation, and 114 out of 109089 from the test set.
3.2 Evaluation Protocols

### v2

@@ Abstract
Accurate automatic screening of minors in unconstrained images requires models robust to distribution shift and resilient to the under-representation of children in public datasets. To address these issues, we propose a multi-task architecture with dedicated under/over-age discrimination tasks based on a frozen FaRL vision-language backbone joined with a compact two-layer MLP that shares features across one age-regression head and four binary underage heads (12, 15, 18, and 21 years). This design focuses on the legally critical age range while keeping the backbone frozen. Class imbalance is mitigated through an $\alpha$-reweighted focal loss and age-balanced mini-batch sampling, while an age gap removes ambiguous samples near thresholds.
Evaluation is conducted on our new Overall Underage Benchmark (303k cleaned training images, 110k test images), defining both the “ASORES-39k” restricted overall test, which removes the noisiest domains, and the age estimation wild-shifts test “ASWIFT-20k” of 20k-images, stressing extreme poses (>>45°), expressions, and low image quality to emulate real-world shifts.
Trained on the cleaned overall set with resampling and age gap, our multiage model “F” reduces the mean absolute error on ASORES-39k from 4.175 y (age-only baseline) to 4.068 y and improves under-18 detection from F2 score of $0.801$ to $0.857$ at $1\%$ false-adult rate. Under the ASWIFT-20k, the same configuration nearly sustains $0.99$ recall while F2 rises from $0.742$ to $0.833$, demonstrating robustness to domain shift.
@@
Age Estimation , Underage Detection , Distribution Shift, Class imbalance
@@ 1 Introduction
Facial age estimation aims to predict a person’s age from a facial image [Angulu_2018_survey]. Although it is a regression task, it is often formulated as classification, by rounding ages or defining age groups [Eidinger_2014_Adience]. The binary case, distinguishing subjects below or above a threshold, is known as age verification or, conversely, as underage or minor detection.
@@
Age estimation and underage detection are increasingly important across a wide range of applications, including age-restricted services, targeted advertising, social media moderation, and, most critically, the protection of minors. Applications involve equipping devices such as gambling or cigarette vending machines, or age-restricted smartphone applications with automated facial age-verification systems. Minor detection is essential in enforcing age-related regulations in digital environments and to combat the distribution of Child Sexual Exploitation Material (CSEM). CSEM hidden in huge amounts of data is detected in a two-step process, consisting of the detection of sexual activity [Gangwar_2024_DeepHSAR] and the detection of underage subjects.
For CSEM detection, the costs of classification errors are asymmetric: a false adult detection means an undetected case of child abuse, while a false minor detection only means that the officer who inspects the detections will find out it is a false detection. Therefore, such systems are typically operated at high recall values, (e.g., 0.99)— to avoid missing minors — and at lower precision (accepting some false alarms).
@@
While significant progress has been made in facial age estimation, most existing solutions are designed for controlled settings and tend to perform well when image conditions are ideal. However, real-world scenarios (“in the wild”) present substantial challenges, including variations in head pose, facial expressions, lighting conditions, image resolution, and frequent occlusions [Jeuland_2022_occlusion]. These factors severely impact the reliability of the model and often degrade its performance. These challenges become even more pronounced when distinguishing between children, adolescents, and adults. This is precisely where accurate detection becomes more critical for legal and safety considerations. Furthermore, publicly available facial datasets, like MORPH [Ricanek_2006_MORPH], AFAD [Niu_2016_AFAD] or AgeDB [Moschoglou_2017_AgeDB], have skewed distributions with a significant under-representation of children and adolescents. This imbalance leads to lower performance for younger age groups. Consequently, the effectiveness of underage detection systems is compromised in critical applications, such as CSEM investigations. In this work, we address the challenges of robust facial age estimation and underage detection with three objectives:
  • A benchmark for age-estimation models under unconstrained conditions.
  • Improving age-estimation models, with focus on such unconstrained scenarios.
  • Enhancing age estimation for minors and the discrimination of minors and adults.
@@
To achieve these, we propose a comprehensive framework for unconstrained environments. The main contributions of this study are as follows:
@@
  • The Overall Underage benchmark and its restricted subset ASORES-39k, designed for the accurate evaluation of age estimation and underage detection models while excluding examples with noisy or unreliable labels.
  • ASWIFT-20k, a novel wild test benchmark, for robustness evaluation in diverse and challenging imaging conditions, including variations in pose, facial expression, illumination, and image quality.
@@
  • An age-balanced sampling strategy during training to mitigate class imbalance. This ensures equal representation of minors, enhancing the ability of the model to generalize across underrepresented age groups.
  • We also explore $\alpha$-reweighted focal loss to mitigate the effects of class imbalance and to prioritize difficult and minority-class samples
The proposed methodology aims to improve facial age estimation, particularly to distinguish between prepubescent, pubescent, and adult individuals. Based on our experience working with law enforcement agencies (LEAs) on previous projects and given the same age for males and females, we established the relevant ages as 12 (prepubescent), 15 (pubescent), and 18 or 21 (adult) [Blanchard2009]. The focus is on developing robust underage detection systems that perform well across varied real-world conditions, including diverse poses, facial expressions, and poor image quality. In addition, it addresses key machine-learning challenges, such as class imbalance and domain shift.
The rest of the paper is organized as follows. Section 2 reviews related work. Sections 3 and 4 describe the proposed methodology (dataset, model architecture and training). Section 5 presents experimental results and discusses performance across different benchmarks. Finally, Section 6 summarizes the main conclusions.
@@ 2 Related Work
Facial age estimation
Facial aging involves gradual changes in facial proportions, feature shape, and skin texture. Early works explicitly modeled these factors. Lanitis_2002_FGNET built a statistical model using principal component analysis of facial landmarks and the isolation of the aging axis in PCA space, while Han_2013_AgeEstimation introduced biologically inspired features with a decision tree and support-vector regressors.
Later approaches employed deep learning and convolutional neural networks (CNNs) [LeCun2015_deeplearning], trained end-to-end without the need for landmarks or handcrafted features [Levi_2015_AgeCNN, Rothe_2018_IMDBWIKI]. Model sizes range from a few million parameters [Levi_2015_AgeCNN] to over 100M parameters in the case of the VGG-16 architecture used by Rothe_2018_IMDBWIKI. SSR-Net [Yang_2018_SSRNet] is a CNN-based age-regression model that uses a stage-wise strategy, reducing the size (less than 100k parameters) and computational effort significantly. Li_2022_CVPR developed an adaptive label distribution learning, exploiting the ordinal nature of age classes.
@@
Recently, Paplham_CVPR2024_benchmark compared state-of-the-art age estimation algorithms and provided an age estimation benchmark, including train/validation/test splits for several facial age estimation datasets. Considering the mean absolute error averaged over the test set of the respective split, they found that the key to good age-estimation accuracy is a large pre-trained backbone rather than the details of the loss function. The proposed model of Paplham_CVPR2024_benchmark combines the FaRL image encoder [Zheng_2022_FaRL] with a small multilayer perceptron (MLP) for age estimation. FaRL is a general facial representation model based on the general-purpose CLIP model [Radford_2021_CLIP], pre-trained with contrastive loss to learn the high-level semantic meaning from image-text pairs for a general facial representation.
Facial underage detection
@@
Meanwhile, other works have focused on the binary task of distinguishing minors from adults. CastrillonSantana_2018 composed a dataset of unrestricted images from a collection of datasets and compared local descriptors and CNNs, i.e., features learned end-to-end, for underage detection. They found that CNNs slightly outperformed local descriptors, but that score fusion among the two approaches could lead to even better accuracy. Anda_2020_DeepUAge proposed the DeepUAge age estimation model, trained with the VisAGe dataset for accurate age estimation of underage subjects, and Chavez_VISAPP2020 studied occluded-eye cases relevant to CSEM. More recently, Gangwar_2021_AttMCNN employed CNNs with an attention mechanism and metric learning for pornographic content detection and age-group classification, and combined these into a CSEM detection pipeline.
3 Overall Underage Benchmark and ASWIFT-20K wild test
@@
This section introduces the proposed dataset and benchmarks: the Overall Underage Benchmark (Section 3.1), together with the ASORES-39k test and the ASWIFT-20K test, designed to assess model robustness to distribution shift (Section 3.2).
3.1 Overall Underage Benchmark
@@
For real-world applications of facial age estimation and underage detection, the age algorithm must be reliable across diverse data and generalize well; robustness to distribution shifts (e.g., imaging conditions, subject demographics, poses, facial expressions) is crucial. Therefore, we merge several benchmarks across the train/validation/test splits provided by Paplham_CVPR2024_benchmark into an overall benchmark. The joint training set consists of data from the publicly available datasets AFAD [Niu_2016_AFAD], AgeDB [Moschoglou_2017_AgeDB], CLAP2016 [Agustsson_2017_AppaReal], the Cross-Age Celebrity Dataset (CACD2000) [Chen_2014_CACD2000], MORPH Album 2 [Ricanek_2006_MORPH], and UTKFace [Zhang_2017_UTKFace]. These datasets cover variations in age, image quality, pose, and facial expression. Figure 1 shows some representative examples.
@@
Figure 1: Sample face patches from the different age estimation benchmarks, as processed by the model during training, i.e., with the random flipping, scaling, and cropping described in Section 4.
@@
AFAD contains Asian subjects with low resolution, and the face crops are tight, which leads to padding in the face patches. MORPH Album 2 contains mugshot images with mostly frontal poses, neutral to negative facial expressions and controlled illumination. AgeDB, CACD2000, CLAP2016, and UTKFace are unconstrained datasets with diversity in pose, illumination, and expression. AgeDB also contains monochrome photographs. Among them, the CACD2000 dataset [Chen_2014_CACD2000] stands out in the sense that it has noisy age annotations and was not initially intended for age estimation. Therefore, we analyze two benchmark variants, with and without the CACD2000 dataset.
Figure 2: The age distribution of the training and test data. (a) Number of images for each source dataset (colors) and for the merged corpus (black). (b) same curves for the whole age range normalized by the respective dataset sizes. (c) Cumulative histogram of the overall test, the ASORES-39k restricted test, and the ASWIFT-20k test and its subsets. Note that the “Pose-45” subset, i.e., those faces with a pose angle beyond 45°, comes mostly from the Dartmouth Dataset of Children’s Faces and thus contains mostly children.
@@
Figure 2 shows the age distribution of the overall training set, resolved by source datasets. Images of subjects below 15 years are only weakly represented, and come mostly from the datasets UTKFace, CLAP2016, and AgeDB. CACD2000 and AFAD, which are concentrated in the age ranges 15 – 35 and 20 – 45, respectively, show a steep rise beyond 15 years. The resulting aggregate displays the highly skewed age distribution that motivates the age-balanced resampling scheme, below in Section 4.2.
@@
We employ the data splits provided by Paplham_CVPR2024_benchmark and use the respective holdout parts of the training data for testing. Further, we use the FG-NET dataset [Lanitis_2002_FGNET] for testing only, as well as the Dartmouth Database of Children’s Faces [Dalrymple_2013_Dartmouth] and the CASIA-Face-Africa dataset collected by the Chinese Academy of Science’s Institute of Automation (CASIA) [Muhammad_2021_Casia].
@@
We used a preliminary model for detecting inconsistencies in the age labels. We selected the images with the largest training and validation errors [^1] Criteria: (i) worst over-18 classifications. labeled age $\geq 18$: lowest $0.5\%$ of over-18 scores ($1738$ images); (ii) worst under-18 classifications. labeled age <18<18: highest $2\%$ of over-18 scores ($293$ images); (iii) worst age underestimations. estimation error $\leq-16$ and estimate <24<24 ($1006$ images); (iv) worst age over-estimations. estimation error $\geq 9$) and labeled age <18<18 ($990$ images). for manual inspection and classified them into the categories bad age, doubtful, fine and bad image. To avoid bias toward our algorithm, doubtful cases were preserved, and only images with clearly wrong age labels and invalid images (non-faces or cartoons) were removed. In another iteration, we also cleaned the test set (with slightly more tolerant settings). In total, 2777 out of 303075 training images, 238 out of 60220 from validation images, and 114 out of 109089 test images were removed.
3.2 ASWIFT-20k, ASORES-39k and Evaluation Protocols